OpenStudy (anonymous):

The Pharaoh Chewing Gum Company has decided to sponsor an additional prize in the radio station's contest. They are giving away backstage passes for the concert! Pharaoh Chewing Gum manufactures a new product they are trying to promote. The new product is a pyramid-shaped gum with a square base. In the spirit of the other challenges, the company has decided to place their pyramid-shaped gum inside a clear glass giant bubble-gum shaped sphere. Each piece of gum has a base

OpenStudy (anonymous):

I would really appreciate help please

OpenStudy (anonymous):

@Riddellikins can I please have some help

OpenStudy (anonymous):

I don't understand the question. What answer are you looking for?

OpenStudy (anonymous):

exactly..what is the question ?

OpenStudy (anonymous):

The Pharaoh Chewing Gum Company has decided to sponsor an additional prize in the radio station's contest. They are giving away backstage passes for the concert! Pharaoh Chewing Gum manufactures a new product they are trying to promote. The new product is a pyramid-shaped gum with a square base. In the spirit of the other challenges, the company has decided to place their pyramid-shaped gum inside a clear glass giant bubble-gum shaped sphere. Each piece of gum has a base measurement of 1 inch and a height of 0.75 inches. The glass sphere container has a diameter of 17.25 inches. How many pieces of Pharaoh Chewing Gum can fit inside the glass sphere? You must show all work to receive credit.

OpenStudy (anonymous):

here is the full question....my bad

OpenStudy (anonymous):

Haha, that's better. Thanks. I will have a look

OpenStudy (anonymous):

To do this roughly, one would get the volume of the sphere and divide it by the volume of the little gum pieces. To do it exactly would be very difficult, as fitting the pieces into the sphere will leave various amounts of open volume, depending on details of the geometry.

OpenStudy (anonymous):

can you please help me do this roughly

OpenStudy (anonymous):

The only way that I know how to do this just by using pure volume measurements. This will give a wrong answer but it may be accurate enough for the question. (as douglaswinslowcooper has said) So you need to find the volume of the sphere and the area of the pyramid and then divide the first by the last: Vsphere = (4/3)Pi r^3 V = (4/3)Pi * 17.25 V = 21500.9 inches cubed Vpyramid = (1/3)w^2*h V = (1/3) * 1*1 * 0.75 V = 0.25 inches cubed 21500.9/0.25 = 86003.6 Therefore it will fit approximately 86003 pieces of gum

OpenStudy (anonymous):

Yes, use the formula for the volume of a sphere, V = (4/3) pi r^3 to get V. Use the formula for the volume of a pyramid with a square base, v = (1/3)(area of base)(height) Estimate maximum number possible in sphere as N = V/v.